A heat pump operates between reservoirs at 299 K (inside the house) and 257 K (outside). It is not a reversible heat pump; it only pumps 74% of the maximum possible heat into the house. How much work must be done to “pump” 6500 J of heat energy into the house?
Answer #1

Here,

to pump 6500 J of heat

ideal heat needed, Q = 6500/0.74 = 8784 J

Now, for the work done W

W/Qinput = (Th - Tc)/Th

W/8784 = (299 - 257)/299

W = 1234 J

the work done needed is 1234 J